OpenStudy (anonymous):

Suppose that four guests check their hats when they arrive at a restaurant, and that these hats are returned to them in a random order when they leave. Determine the probability that no guest will receive the proper hat.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i think we can do this

OpenStudy (anonymous):

the probability that the first one does not get her hat is \(\frac{3}{4}\)

OpenStudy (anonymous):

leaving 3 hats, so the probability that the second guest does not get her hat is \(\frac{2}{3}\)

OpenStudy (anonymous):

hmm maybe that is not right, hold the phone

OpenStudy (anonymous):

definitely wrong, sorry lets try again

OpenStudy (anonymous):

mb we should try with 3 persons to ease?

OpenStudy (anonymous):

to find probability of 3 people not getting their hat instead of 4

OpenStudy (anonymous):

this is called a "derangement" there is a formula for it

OpenStudy (anonymous):

or if you don't want to use the formula, count there are \(4!\) ways to distribute the 4 hats of those i believe there are 9 ways that they do not match \(4!=24\) so it is not too many to list

OpenStudy (anonymous):

i.e. start with ABCD as the correct order, then start listing ABCD (4 matches) ABDC (two matches) ACBD (2 match) ACDB (1 match) ADBC (1 match) ADCB (2 match) obviously you have to start with something other than A

OpenStudy (anonymous):

if i am not mistaken you will count 9 and get \(\frac{9}{24}\) but we can also try the formula \[\sum_{k=0}^4\frac{(-1)^k}{k!}\] and see if we get the same answer

OpenStudy (anonymous):

wow how to you like that! http://www.wolframalpha.com/input/?i=sum+k+%3D0+to+4+%28-1%29^k%2Fk!

OpenStudy (anonymous):

wait a minute i'll try to count it manually

OpenStudy (anonymous):

i got 7/24

OpenStudy (anonymous):

thank you!